The same enhancement happened around 01:20 UTC on September 27th. The enhancement feature, duration (~ 20 minutes) are almost the same with the recent two events.
Also ~ 5K level temporal enhancement at the 1st head of the cryocooler for the BRT side happened in EXC about 53 days ago from now.
They fortunately recovered.
According to Kimura-san, any detected errors result in the stop of the cryocooler and no automatic restart even when the error sources are recovered. The water supply system shows no errors according to the remote webcam image. So, water supply system seems to be healthy.
In this time in IXC, the warming up time by 15K took 20 minutes. While, it took a shorter time of 15 minutes when the cryocoolers stopped because of the power outage in the EX area on October 28th. So in IXC, the cryocooler itself kept working, but lost its cooling power because of something.
The same warm-up happened again on Sunday mid-night.
The temp change of the 1st cryocooler head was a very smooth enhancement like a natural warmup. Did a temporal stop of something happen?
[Komori, KYamamoto]
We started the frequency scan to identify IY resonant frequencies.
We found several resonant frequencies as follows:
- 15.795 kHz
- 16.1928 kHz
- 16.2536 kHz
- 21.0635 kHz
- 21.0757 kHz
[Komori, KYamamoto]
We performed the EX Q-measurement as follows:
Resonant frequency: 23.9261 kHz
Q: (2.41 ± 0.13) x 10^5
This is the final Q-measurement for EX low frequencies.
From this measurement, I began using the improved Python code for Q-measurement (/users/Commissioning/Q_meas/Qmeasfit.py).
The main difference is a reduction in the number of fitted parameters.
In the previous code, the excitation period was included in the fitting, and the moment when the ringdown starts was one of the fitted parameters.
This setup led to frequent fitting errors due to the fluctuating excited amplitude.
Therefore, I excluded the excitation period from the fitting region and removed the moment when the ringdown starts from the fitted parameters.
The results of the investigation of scattered light noise conducted on 12/08 (klog#27877) are summarized in a slide.
Please see JGWdoc.
[Komori, KYamamoto]
We measured the Q-value of an EX mechanical mode and reconfirmed the resonant frequency and Q-value of the fundamental drumhead mode.
The results, based on 10 measurements, are as follows:
- Resonant frequency: 21.07462 kHz
- Q: (1.84 ± 0.23) x 10^5
- Resonant frequency: 23.6928 kHz
- Q: (1.07 ± 0.04) x 10^5
For the time being, please use this frequency as the beacon sideband.
This frequency differs by 2.6 Hz from the previous report (klog:27634).
This discrepancy might be due to the fact that the peak of the transfer function does not precisely indicate the resonant frequency because of beating with the injection frequency.
However, the actual resonant frequency might indeed be changing, as reported in klog:27893.
From this measurement, I began using the improved Python code for Q-measurement (/users/Commissioning/Q_meas/Qmeasfit.py).
The main difference is a reduction in the number of fitted parameters.
In the previous code, the excitation period was included in the fitting, and the moment when the ringdown starts was one of the fitted parameters.
This setup led to frequent fitting errors due to the fluctuating excited amplitude.
Therefore, I excluded the excitation period from the fitting region and removed the moment when the ringdown starts from the fitted parameters.
Figure 1 shows the generated C code of a part of k1visetmxp before (left) and after (right) fixing the problem.
As shown in the highlight line, the input to the 16th bit of bit2word was "sum4" instead of a proper input signal. "sum4" is a summation of yaw contorl signals from ASC, ADS, and BPC. And the 16th bit is assigned to switch the coil driver input as Drive-In or Test-In for the control signal on 4-9 pin. So the model-compile script generated a very nice code which disabled the output to one coil when yaw control signal became non-zero value.
I tried to modify the conversion script from the matlab code to the C code as shown below. After this modification, real-time models were re-compiled and worked fine.
--- Bit2Word.pm.bak 2016-11-09 09:33:17.000000000 +0900
+++ Bit2Word.pm 2023-12-09 18:25:19.000000000 +0900
@@ -75,7 +75,7 @@ sub frontEndCode {
for (0 .. $::partInCnt[$i]-2) {
$calcExp .= "\t$::fromExp[$_],\n";
}
- $calcExp .= "\t$::fromExp[$partInCnt[$i]-1]\n";
+ $calcExp .= "\t$::fromExp[$::partInCnt[$i]-1]\n";
$calcExp .= "};\n";
$calcExp .= "\L$::xpartName[$i] = 0;\n";
$calcExp .= "for (ii = 0; ii < $::partInCnt[$i]; ii++)\n{\n";
ETMY suddenly tripped (went to PAY_TRIPPED) at 19:46:28 JST because IM_H3 sensor detected larger RMS than the watchdog threshold as shown in Fig.1. V3 and H2 also detected similar behavior but they were much smaller than the threshold.
On the MN stage, no strange behavior was detected on local damp control signals before ETMY was tripped. Large changes in Fig.2 came from the state transition from LOCK_ACQUISITION to TRIPPED. So it shouldn't be caused by local control. I also couldn't find any coincident behavior on PEM channels around EY. The behavior detected on local sensors is so fast for a pendulum motion that it may be an electrical issue.
Because there was no strange behavior on IM local sensors after ETMY was tripped, I cleared the DACKILL and watchdogs and recovered LOCK_AQUISITION state.
[Yamamoto, Ushiba]
We found the exact reason why PRFPMI cannnot be locked with ASC: it is not due to CPU load but the problem in C code of the real-time model.
Somehow, ISC_Y signals send to 16th bit of BO instead of real BO signal and change the coil driver setting when VIS model obtained ISC_Y signals from ASC, BPC, and ADS models.
So, we could not use any alignment controls with ISC_Y signals including initial alignment, and please do not use ISC_Y controls until fixing the bug.
To reduce filter banks further, I deleted the following filter banks:
K1:VIS-{ITMX,ITMY,ETMX,ETMY}_TM_ADD_{L,P,Y}
Figure 1 and 2 show the part of models before and after modification.
During the work, I restarted DAQ at around 16:00.
To reduce the CPU load, I cleaned up filter baks of Type-A payload.
So, please see the archive at 2023/12/09 00:00:00 UTC, if you want to see the filters befroe the clean up.
The 1st head temp of the cryocooler for the BS side duct shield in IXC showed ~15K enhancement and recovery around the early morning today.
The vacuum level at BS seems to show no remarkable enhancement.
Temp enhancement in the BS side duct shield seemed to be within ~1K around 90K~100K.
[kTanaka, Hirose]
We checked f3 signal to go to EOM2b. We found that the f3 signal was coming to the input of the EOM with the same modulation index as f1. We noticed that the N-SNA conversion adapter was loose during the process. We would like to measure the modulation depth again in this condition.
## Detail
## Process
This time we checked the RFamp input and output, taking care not to unplug the cable while the signal was going from Seed's function generator. We summarized for future use.
Maybe due to the real-time model update (klog27894), CPU over load occured when we engaged {D,C}HARD_Y loops and PRFPMI lost lock.
Now, we cannot reach PRFPMI_RF_LOCKED state, so please do not request PRFPMI_RF_LOCKED state to LSC_LOCK guardian until the issue will be solved.
I will try to fix the issue but If you want to use IFO before that, please use PRFPMI_TRANSITION_TO_1FSIG state.
[Yamamoto, Ushiba]
We found the exact reason why PRFPMI cannnot be locked with ASC: it is not due to CPU load but the problem in C code of the real-time model.
Somehow, ISC_Y signals send to 16th bit of BO instead of real BO signal and change the coil driver setting when VIS model obtained ISC_Y signals from ASC, BPC, and ADS models.
So, we could not use any alignment controls with ISC_Y signals including initial alignment, and please do not use ISC_Y controls until fixing the bug.
Figure 1 shows the generated C code of a part of k1visetmxp before (left) and after (right) fixing the problem.
As shown in the highlight line, the input to the 16th bit of bit2word was "sum4" instead of a proper input signal. "sum4" is a summation of yaw contorl signals from ASC, ADS, and BPC. And the 16th bit is assigned to switch the coil driver input as Drive-In or Test-In for the control signal on 4-9 pin. So the model-compile script generated a very nice code which disabled the output to one coil when yaw control signal became non-zero value.
I tried to modify the conversion script from the matlab code to the C code as shown below. After this modification, real-time models were re-compiled and worked fine.
--- Bit2Word.pm.bak 2016-11-09 09:33:17.000000000 +0900
+++ Bit2Word.pm 2023-12-09 18:25:19.000000000 +0900
@@ -75,7 +75,7 @@ sub frontEndCode {
for (0 .. $::partInCnt[$i]-2) {
$calcExp .= "\t$::fromExp[$_],\n";
}
- $calcExp .= "\t$::fromExp[$partInCnt[$i]-1]\n";
+ $calcExp .= "\t$::fromExp[$::partInCnt[$i]-1]\n";
$calcExp .= "};\n";
$calcExp .= "\L$::xpartName[$i] = 0;\n";
$calcExp .= "for (ii = 0; ii < $::partInCnt[$i]; ii++)\n{\n";
We tried measuring the Q-value of a mode that we couldn't measure previously, and we succeeded in doing so for one mode.
This records the highest mirror Q-value ever in KAGRA.
Resonant frequency: 23.92353 kHz
Q: 4.3x10^6 and 4.0x10^6
We made 5 attempts, but the interferometer was down for the last 3 measurements, so we only succeeded in the first 2 attempts.
Interestingly, the resonant frequency differed by 0.14 Hz compared to that of yesterday, while the mirror temperature remained almost the same (0.02 K defference).
Particularly from the viewpoint of the beacon frequency, we should keep this in mind.
We finished the Q-measurement of IX mechanical modes below 30 kHz.
Ushiba-san, ikeda
Glitch was caused by a change in K-Log#27894.
Therefore, some filter banks of Type-A were deleted.
Deleted:
K1:VIS-{ETMX,ETMY,ITMX,ITMY}_IM_OPTICALIGN
K1:VIS-{ETMX,ETMY,ITMX,ITMY}_IM_PSD
K1:VIS-{ETMX,ETMY,ITMX,ITMY}_TM_OPTICALIGN
I removed the GigE camnera at the PR2 and used another camera ASI294MC which is being used as Tcam and has a good sensitivity for IR.
I could see the IR spot. The white area in Fig.1 circled by Magenta lines comes from IR light since RGB has the same response above IR wavelength. Today, all the suspensions are in safe mode, so the IR spot was hitting somewhere from the center of the mirror.
The purpose of this new camera is to see the scattering IR light clipped by baffles or something during lock. In this meaning, This camera meets the requirements.
I saw a lot of red light from Oplev, but I did not see green light coming from the green laser. Even direct watching by my eye, I saw only red light and I did not see the green light at all.
The focal length of this lens was a bit too long, so the area we can see is a bit small. Comparing Fig.2 and Fig.3 can help to identify where it is. Both are exactly the same angle.
Today's camera and lens are just a test version, I need to prepare another filter or shorter focal length lens. I will order a proper camera and lens and some adapters. I put the GigE camera back today. When the new camera is ready, I will install it.
[Yamaguchi, Nakagaki]
We installed Wi-Fi access points in the Xend and Yend machine rooms.
The SSIDs are
EX-MachineRoom and
EY-MachineRoom.
We have identified additional resonant frequencies for ETMX through differential excitation as follows:
We could not detect mechanical modes around 27 kHz.
We have completed the scanning for EX low resonances.
kTanaka-san, ikeda
Continuing from last week, we installed PMC pico motor cables in the PSL room.
From the PSL room mesh shelves to the back side of the PSL room.
It is not yet connected to a motor driver.
It will be connected at the next PMC mount replacement work.
start time: 2023/12/08 10:34:50(JST), 1386034508
End Time: 2023/12/08 11:12:03(JST), 1386036741